OpenStudy (anonymous):

Choose the Standard Form of the equation y = - x - 6.

OpenStudy (ash2326):

if it's a line then it's in standard form y=-x-6 other wise x+y=-6

OpenStudy (anonymous):

But they give you these choices 1/4 x + y = -6 4x + y = -6 x + 4y = -24 -4x + y = -24

OpenStudy (ash2326):

i think you need to check the question??

OpenStudy (anonymous):

Choose the Standard Form of the equation y = - x - 6. Answer 1/4 x + y = -6 4x + y = -6 x + 4y = -24 -4x + y = -24

OpenStudy (ash2326):

you need to have a term with 4 in the question??

OpenStudy (anonymous):

Jeezy oh petes, I don't quite know that's why i'm here LOL

OpenStudy (ash2326):

Brookeluda if we were given y=-x/4-6 i could write it like 4y+x=-24 but we can't guess. You check what's the equation from your textbook or paper

OpenStudy (anonymous):

Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (0, 4) and (-2, 8) (3, 7) and (-1, -5) Answer The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. The slope of Set #1 is smaller than the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

Could ya help me with that! ^^^